OpenStudy (anonymous):

will give medal and fan! help please?(:

OpenStudy (anonymous):

@marybel86

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

can u find slope given 2 points ?

hartnn (hartnn):

\(\large m= \dfrac{y_2-y_1}{x_2-x_1}\) where, x1,y1 = 1,3 x2,y2 =3,17 just plug in values!

OpenStudy (anonymous):

Oh so 17- 3/ 3-1 ? @hartnn

hartnn (hartnn):

yes! so, m=... ?

hartnn (hartnn):

whats 17-3=... ?

OpenStudy (anonymous):

14 :)

OpenStudy (anonymous):

do you now do 14/2 ? @hartnn

hartnn (hartnn):

so, m = 14/2 =... ?

OpenStudy (anonymous):

7 ? (:

OpenStudy (anonymous):

aren't the lines perpendicular?

OpenStudy (anonymous):

yup!

jimthompson5910 (jim_thompson5910):

so that would mean 7 isn't your final answer, but it looks like you know what you're doing

OpenStudy (anonymous):

wait so then do i sq root it?

jimthompson5910 (jim_thompson5910):

if m is the slope of one line then -1/m is the slope of the line perpendicular to that first line

jimthompson5910 (jim_thompson5910):

ex: if 2/3 is the slope of one line, you flip the fraction and the sign to get -3/2 So -3/2 is the slope of the perpendicular line (it's perpendicular to the line with a slope of 2/3)

OpenStudy (anonymous):

\[m_1*m_2=-1\]

OpenStudy (anonymous):

oh so its like - 2/14 instead? (:

jimthompson5910 (jim_thompson5910):

which reduces to ???

OpenStudy (anonymous):

negative reciprocal :) found it!

hartnn (hartnn):

sorry. i lost connection :(

OpenStudy (anonymous):

-0.14

jimthompson5910 (jim_thompson5910):

keep it as a fraction though

jimthompson5910 (jim_thompson5910):

reduce it as much as possibl

jimthompson5910 (jim_thompson5910):

possible*

OpenStudy (anonymous):

what's - 0.14 as a fraction? (:

hartnn (hartnn):

reduce -2/14 how many times 2 goes into 14 ?

OpenStudy (anonymous):

14 times?

hartnn (hartnn):

really ? i thought 7 times..

hartnn (hartnn):